### v1

@@ Abstract
Quantum adiabatic optimization has long been expected to outperform classical methods in solving NP-type problems. While this has been proven in certain experiments, its main applications still reside in academic problems where the size of the system to be solved would not represent an obstacle to any modern desktop computer. Here we develop a systematic procedure to find the global optima of the modularity function to discover community structure in complex networks solely relying on pure annealers rather than hybrid solutions. We bypass the one-hot encoding constraints by hierarchically and recursively encoding binary instances of the problem that can be solved without the need to guess the exact penalties for the Lagrange multipliers. We study the variability, and robustness of the annealer as a function of network size, directness of connections, topology, and the resolution of the communities. We show how our approach produces meaningful and at least equally optimal solutions to state-of-the-art community detection algorithms while maintaining tractable computing times. Lastly, due to its recursive nature, the annealing process returns intermediate subdivisions thus offering interpretable rather than black-box solutions. These dendrograms can be used to unveil normal and pathological hidden hierarchies in brain networks hence opening the door to clinical workflows. Overall, this represents a first step towards an applicable practice-oriented usage of pure quantum annealing potentially bridging two segregated communities in modern science and engineering; that of network science and quantum computing.
@@
Quantum Annealing, Modularity Maximization, Community Detection, Complex Networks
@@ Introduction
A graph or network $G=(V,E)$ is a set of vertices or nodes $V$ and edges $E$. Moreover, a community in the graph is a subset of nodes $c\subseteq V$. A community structure is a set of non-overlapping communities $C=\{c_{r}|c_{r}\subseteq V,c_{s}\cap c_{r}=\varnothing\}$. The set of all possible community structures is denoted as $\mathcal{C}$. Importantly, a community is not a subgraph as it doesn’t require the deletion of inter-community edges. Lastly, the quality of a given community structure $C$ can be quantified with the modularity index [7],
@@
A certain community structure is said to be modular if it exhibits a high modularity index when compared to multiple random instances of a given class of random networks (e.g., Erdős-Renyi). Community detection algorithms aim to find a given community structure $C\in\mathcal{C}$ whose associated modularity function $Q$ lies at a global maximum - or its opposite at a minimum,
@@
Even though $\mathcal{C}$ is bounded, the number of possible elements in it grows faster than exponentially with the size of the network [27]. Furthermore, it has been proven that maximizing modularity is an NP-complete problem [17], thus justifying QA methods [18] (or at least inspired). To maximize the modularity index, the first term in Eq. 1 suggests defining clusters containing a large number of edges. Contrary, the second term aims that the size of such clusters should be as small as possible. This is important because any algorithm could be prone to isolate given nodes to artificially boost the modularity index [25].
@@
where $m=\frac{1}{2}\sum_{ij\in G}A_{ij}$ is the total edge-weight, $A_{ij}$ are the entries of the symmetric adjacency matrix,
$B_{ij}=A_{ij}-\gamma\frac{g_{i}g_{j}}{2m}$ (4)
are the entries of the modularity matrix, and $g_{i}=\sum_{j}A_{ij}$ is the node weighted degree. For directed networks [28],
@@
where $m=\sum_{ij}A_{ij}$ is the total edge-weight, $A_{ij}$ are the entries of the asymmetric adjacency matrix,
@@
is the directed modularity matrix, $g_{i}^{in}=\sum_{j}A_{ij}$ is the node weighted in-degree, and $g_{j}^{out}=\sum_{i}A_{ij}$ is the node weighted out-degree. We detailed all the algebraic intermediate steps in the Supplementary Material.
@@
Minimizing $\hat{Q}^{\mathcal{P}}$ need not be straightforward but optimizing for each step of the process will yield the same outcome,
@@
Formally, there is no guarantee that a given optimized process will univocally yield the partition with the highest modularity index in Eq. 2, $\min\left[\hat{Q}^{\mathcal{P}}\right]\to\hat{Q}$. However, given that the problem itself is NP-complete [17], heuristic processes are needed. Several algorithms were developed around the fact that optimizing in smaller steps instead of all at once yields competitive results, the Louvain method being the most popular attempt [10].
Our heuristic is based on the additive nature of $Q$, where each community represents an isolated entity that independently contributes to its value [7]. A natural way to exploit this is to recursively divide communities into two [3], thus,
@@
where $10\leq|C^{k}|-|C^{k-1}|\leq|C^{k-1}|+$ 1, and $Q(C^{k})$ is the modularity of the network after splitting a given community $c\in C^{k-1}$ into two. Thus, we wish to maximize the modularity by recursively subdividing each community into two. It is straightforward to see that $\hat{Q}^{\mathcal{P}}\to Q(C^{\infty})$ with $C^{\infty}$ being a community structure of arbitrary size. In what follows, we derive the Quadratic Unconstrained Binary Optimization (QUBO) cost function for each element of $\mathcal{P}$. We then show how following the approach described here can yield high-quality community structures of arbitrary size without having to use one-hot encoding constraints [23, 24] nor discrete variables [25].
@@
Precisely, the modularity of an undirected network of $|C^{k-1}|$ different communities can be decomposed as
@@
where $k_{1}\cup k_{2}=k$. Subtracting both expressions and using Eq. 3 we obtain a compact expression for the elements of $\mathcal{P}$ in terms of the modularity matrix,
$\begin{split}p_{k}&=\frac{-1}{4m}\sum_{ij\in k}\left(B_{ij}s_{i}s_{j}-B_{ij}\right)-\frac{1-\gamma}{2}\\ &=\frac{-1}{4m}\sum_{ij\in k}\left(B_{ij}-\delta_{ij}\sum_{r\in k}B_{ir}\right)s_{i}s_{j}=\frac{-1}{4m}\sum_{ij\in k}B_{ij}^{\text{k}}s_{i}s_{j}\\ &=\frac{-1}{m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}x_{j}+\frac{1}{2m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}+\frac{1}{2m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{j}+\frac{1}{4m}\sum_{ij\in k}B_{ij}^{\text{k}}\\ &=\frac{-1}{m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}x_{j}\end{split}$ (10)
where we have arbitrarily dropped the constant term $\frac{1-\gamma}{2}$ (i.e., it doesn’t play a role in the optimization), $\delta_{ij}$ is the Kronecker delta function, we used $s_{i}^{2}=1$, and
@@
is the generalized modularity matrix Bk for the $k$-th community [3]. In the second row, we have applied a linear mapping from the Ising variables to a set of binary ones $s_{i}=2x_{i}-1$ to obtain the QUBO formulation. The last row follows from some of the properties of the generalized modularity matrix.
@@
If the network has some degree of asymmetry, we could proceed in an identical manner as for undirected networks. The only difference would be that usage of the expressions for directed networks in order to obtain the elements $p_{k}$ of the process $\mathcal{P}$ (see Supplementary Material for the full derivation).
$p_{k}=\frac{-8}{m}\left[\sum_{ij\in k}B_{ij}^{d,\text{k}}x_{i}x_{j}-\frac{1}{2}\sum_{j\in k}x_{j}\sum_{i\in k}\left(B_{ij}^{d}-B_{ji}^{d}\right)\right]$ (12)
However, as opposed to the undirected case, although $p_{k}$ reduces to the same expression in Eq. 10 if the modularity matrix is symmetric, the first element of the process does not coincide with the modularity function to optimize for the full graph (i.e., $p_{0}\neq Q(C^{0})$). This is undesirable (but not critical) because it would mean that the first step of the hierarchy needs to be separated from the rest. Furthermore, the symbolic expression for the corresponding QUBO problem would possess $N(N+1)$ symbolic terms, potentially compromising the computational workflow for large graphs. Lastly, the $ij$-th and $ji$-th entries would be associated with the product of binary variables $x_{i}x_{j}$. Then, the QPU would interpret the coupling of the two qubits as a sum of $B_{ij}^{d,\text{k}}+B_{ji}^{d,\text{k}}$, hence loosing all the non-symmetric information.
@@
The optimal solution, in terms of the modularity index, the community structure, and the total computational time used, was defined as the solution that yielded the highest modularity index (see Algorithm 2). For that, we ran the hierarchical process 20 times independently. In each one of these runs, the returned modularity was kept only if its value was higher than the previously assigned one (line 7). The D-Wave Advantage sampler was initialized on every one of these runs. The elapsed time was measured from the start of the hierarchical process until the last subdivision was attained; not including the computation of the modularity matrices, the QUBO functions, or the embedding times. It should be noted that the time we report is highly dependent on the specific configuration of the local machine. Therefore, exact quantitative analyses of this variable may not be directly interpretable; nonetheless, the relative times for different runs possess valuable information.
@@
The QUBO problem, illustrated by Eq. 10, can be depicted as a weighted graph. In this representation, the vertices correspond to binary variables $x_{i}$, and the edges signify the quadratic terms $x_{i}x_{j}$ in the problem formulation. Linear coefficients are attributed as weights to the vertices, while quadratic coefficients are similarly assigned to the edges.
The actual D-Wave QPU topology uses a Pegasus graph of degree 15 [29]. The QUBO graph must be mapped onto that architecture, which is achieved by minor-embedding [30] i.e. mapping each logical binary variable to a group of physical qubits on the actual machine so all required connections are realized. As the QUBO function based on generalized modularity matrix (see Eq. 10) is dense, we used complete graph embedding predefined routine (clique embedding). To reduce the time of the embedding procedure, we used the cache mechanism to compute embedding for all necessary clique sizes only once and reuse it while performing hierarchical approach described by Algorithm 1.
@@
To compare the performance of the hierarchical process and pure quantum annealing, we evaluated results using four different approaches: the Louvain, Leiden, Bayan, and Hierarchical Gurobi algorithms, as well as a hybrid quantum-classical method based on discrete quadratic optimization models (DQM) [10, 12, 14, 25]. For the Leiden and Louvain algorithms, we computed multiple instances for each network. Specifically, each algorithm was run a certain number of times per network, and the community structure with the highest modularity was retained as the ”true” structure. The DQM method demonstrated high stability, likely due to its problem partitioning strategy for fitting into the annealer [24, 25]. The Hierarchical Gurobi used the same algorithms as the Hierarchical annealing but only used the classical Gurobi solver. For the Bayan, which relies heavily on the Gurobi optimizer and its corresponding license, the Louvain, and Leiden algorithms we ran 20 instances per network. The Bayan algorithm is known to return exact solutions for networks up to a considerable size [15].
Community detection algorithms often yield varying structures due to the use of heuristics or inherent solution degeneracy [31]. We evaluated the variability between communities identified by two algorithms through: 1) consensus classification matrices and 2) the Dice-Sørensen score. For each network, an $N\times N$ consensus coclassification matrix recorded how often nodes $i$ and $j$ were grouped into the same community across algorithm runs [32]. Perfect agreement between algorithms would result in matrix entries $M_{ij}$ of 0 or 1 (or 0% and 100%), indicating stable or well-defined community structures. Conversely, a range of intermediate $M_{ij}$ values suggests inconsistent community assignments. While informative, this matrix does not assess community structure optimality, necessitating the additional use of a modularity index. To compare community structures directly, we computed the Dice score between pairs of communities as:
@@
where $Z$ is in-place for any of the classical alternatives described.
@@
We generated random networks using three different models. First, the Erdős-Rényi model [33] connected any pair of initially disconnected nodes with a constant probability $p$, producing random networks without clear statistical organization. Despite this, non-trivial geometrical behaviors can still emerge, making it a useful null model for study. Second, the Barabasi-Albert model [34] added new nodes dynamically, where the probability of a new node $i$ connecting to node $j$ was proportional to $j$’s degree, $p_{j}\propto k_{j}$. This preferential attachment mechanism produced networks with robust scaling properties and enhanced resilience. Finally, we used the Holme-Kim model [35], which incorporates preferential attachment but adds a triad formation step to boost clustering, defined as the presence of connected triangles in the network. This produced highly clustered, power-law networks.
For these networks, we fitted a discrete power-law distribution to their node degrees present in a given network [36]. A network is said to have a power-law distribution or structure if the probability $P(k)$ of a given node to have degree $k$ follows an exponential distribution,

### v2

@@ Abstract
Quantum adiabatic optimization has long been expected to outperform classical methods in solving NP-type problems. While this has been proven in certain experiments, its main applications still reside in academic problems where the size of the system to be solved would not represent an obstacle to any modern desktop computer. Here we develop a systematic procedure to find the global optima of the modularity function to discover community structure in complex networks solely relying on pure annealers rather than hybrid solutions. We bypass the one-hot encoding constraints by hierarchically and recursively encoding binary instances of the problem that can be solved without the need to guess the exact penalties for the Lagrange multipliers. We study the variability, and robustness of the annealing process as a function of network size, directness of connections, topology, and the resolution of the communities. We show how our approach produces meaningful and at least equally optimal solutions to state-of-the-art community detection algorithms while maintaining tractable computing times. Lastly, due to its recursive nature, the annealing process returns intermediate subdivisions thus offering interpretable rather than black-box solutions. These dendrograms can be used to unveil normal and pathological hidden hierarchies in brain networks hence opening the door to clinical workflows. Overall, this represents a first step towards an applicable practice-oriented usage of pure quantum annealing potentially bridging two segregated communities in modern science and engineering; that of network science and quantum computing.
@@
Quantum Annealing, Modularity Maximization, Community Detection, Complex Networks, Brain Connectivity
@@ Introduction
A graph or network $G=(V,E)$ is a set of vertices or nodes $V$ and edges $E$. Moreover, a community in the graph is a subset of nodes $c\subseteq V$. A community structure is a set of non-overlapping communities $C=\{c_{r}|c_{r}\subseteq V,c_{s}\cap c_{r}=\varnothing,\text{ for }s\neq r\}$. The set of all possible community structures is denoted as $\mathcal{C}$. Importantly, a community is not a subgraph as it doesn’t require the deletion of inter-community edges. Lastly, the quality of a given community structure $C$ can be quantified with the modularity index [7],
@@
A certain community structure is said to be modular if it exhibits a high modularity index when compared to multiple random instances of a given class of random networks (e.g., Erdős-Renyi). Community detection algorithms aim to find a given community structure $C\in\mathcal{C}$ whose associated modularity function $Q$ lies at a global maximum – or its opposite at a minimum,
@@
Even though $\mathcal{C}$ is bounded, the number of possible elements it contains grows faster than exponentially with the size of the network [27]. Furthermore, it has been proven that maximizing modularity is an NP-complete problem [17], thus justifying QA methods [18]. To maximize the modularity index, the first term in Eq. 1 suggests defining clusters containing a large number of edges. On the contrary, the second term aims to minimize the size of such clusters. This is important because any algorithm could be prone to isolate given nodes to artificially boost the modularity index [25].
@@
where
$B_{ij}=A_{ij}-\gamma\frac{g_{i}g_{j}}{2m},$ (4)
are the entries of the modularity matrix, $A_{ij}$ are the entries of the symmetric adjacency matrix, $m=\frac{1}{2}\sum_{ij\in G}A_{ij}$ is the total edge-weight, and $g_{i}=\sum_{j}A_{ij}$ is the node weighted degree. For directed networks [28],
@@
where
@@
are the entries of the directed modularity matrix, $A_{ij}$ are the entries of the asymmetric adjacency matrix, $m=\sum_{ij}A_{ij}$ is the total edge-weight, $g_{i}^{in}=\sum_{j}A_{ij}$ is the node weighted in-degree, and $g_{j}^{out}=\sum_{i}A_{ij}$ is the node weighted out-degree. The superscript $d$ denotes a directed or asymmetric matrix. We detailed all the algebraic intermediate steps in the Supplementary Material.
@@
Crucially, minimizing $\hat{Q}^{\mathcal{P}}$ need not be straightforward. However, sequentially optimizing each element of the process $p_{k}$ is entirely equivalent,
@@
Our heuristic is based on the additive nature of $Q$, where each community represents an isolated entity that independently contributes to its value [7]. A natural way to exploit this property is to recursively divide communities into two [3]. Therefore,
@@
where $10\leq|C^{k}|-|C^{k-1}|\leq|C^{k-1}|+$ 1, and $Q(C^{k})$ is the modularity of the network after splitting a given community $c\in C^{k-1}$ into two. It is straightforward to see that $\hat{Q}^{\mathcal{P}}\to Q(C^{\infty})$ with $C^{\infty}$ being a community structure of arbitrary size. Thus, we wish to maximize the modularity by recursively subdividing each community into two.
Formally, there is no guarantee that a given optimized process will univocally yield the partition with the highest modularity index in Eq. 2, $\min\left[\hat{Q}^{\mathcal{P}}\right]\to\hat{Q}$. However, given that the problem itself is NP-complete [17], heuristic processes are needed. Several algorithms have been developed based on the idea that optimizing in smaller steps, rather than all at once, yields competitive results. Among these, the Louvain method is the most popular [10]. In what follows, we derive the Quadratic Unconstrained Binary Optimization (QUBO) cost function for each element of $\mathcal{P}$. We then show how following the approach described here can yield high-quality community structures of arbitrary size without having to use one-hot encoding constraints [23, 24] nor discrete variables [25].
@@
Following Eq. 1, the modularity of an undirected network of $|C^{k-1}|$ different communities can be decomposed as
@@
where $k_{1}\cup k_{2}=k$. Using Eq. 3 we obtain a compact expression for the elements of $\mathcal{P}$ in terms of the modularity matrix,
$\begin{split}p_{k}&=-\left[Q(C^{k})-Q(C^{k-1})\right]\\ &=\frac{-1}{4m}\sum_{ij\in k}\left(B_{ij}s_{i}s_{j}-B_{ij}\right)-\frac{1-\gamma}{2}\\ &=\frac{-1}{4m}\sum_{ij\in k}\left(B_{ij}-\delta_{ij}\sum_{r\in k}B_{ir}\right)s_{i}s_{j}=\frac{-1}{4m}\sum_{ij\in k}B_{ij}^{\text{k}}s_{i}s_{j}\\ &=\frac{-1}{m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}x_{j}+\frac{1}{2m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}+\frac{1}{2m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{j}+\frac{1}{4m}\sum_{ij\in k}B_{ij}^{\text{k}}\\ &=\frac{-1}{m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}x_{j}.\end{split}$ (10)
We have arbitrarily dropped the constant term $\frac{1-\gamma}{2}$ since it doesn’t play a role in the optimization.
@@
is the generalized modularity matrix Bk for the $k$-th community [3], we used $s_{i}^{2}=1$, and $\delta_{ij}$ is the Kronecker delta function. In the second row, we have applied a linear mapping from the Ising variables to a set of binary ones $s_{i}=2x_{i}-1$ to obtain the QUBO formulation. The last row follows from some of the properties of the generalized modularity matrix.
@@
If the network has some degree of asymmetry, we could proceed identically as for undirected networks. The only difference would be the usage of the expressions for directed networks in order to obtain the elements $p_{k}$ of the process $\mathcal{P}$ (see Supplementary Material for the full derivation)
$p_{k}=\frac{-8}{m}\left[\sum_{ij\in k}B_{ij}^{d,\text{k}}x_{i}x_{j}-\frac{1}{2}\sum_{j\in k}x_{j}\sum_{i\in k}\left(B_{ij}^{d}-B_{ji}^{d}\right)\right].$ (12)
However, contraty to the undirected case, although $p_{k}$ reduces to the same expression in Eq. 10 if the modularity matrix is symmetric, the first element of the process does not coincide with the modularity function to optimize for the full graph (i.e., $p_{0}\neq Q(C^{0})$). This is undesirable (but not critical) because it would mean that the first step of the hierarchy needs to be separated from the rest. Furthermore, the symbolic expression for the corresponding QUBO problem would possess $N(N+1)$ symbolic terms, potentially compromising the computational workflow for large graphs. Lastly, the $ij$-th and $ji$-th entries would be associated with the product of binary variables $x_{i}x_{j}$. Then, the QPU would interpret the coupling of the two qubits as a sum of $B_{ij}^{d,\text{k}}+B_{ji}^{d,\text{k}}$, hence loosing all the non-symmetric information.
@@
The optimal solution, in terms of the modularity index, the community structure, and the total computational time used, was defined as the solution that yielded the highest modularity index (see Algorithm 2). For that, we ran the hierarchical process 20 times independently. In each one of these runs, the returned modularity was kept only if its value was higher than the previously assigned one (line 7). The D-Wave Advantage sampler was initialized on every one of these runs. The elapsed time was measured from the start of the hierarchical process until the last subdivision was attained; not including the computation of the modularity matrices, the QUBO functions, or the embedding times. It should be noted that the time we report is highly dependent on the specific configuration of the local machine. Therefore, exact quantitative analyses of this variable may not be directly interpretable; nonetheless, the relative times for different runs possess valuable information. We provide a general description of the software package together with snippets of the high-level Python code developed in the Supplementary Material, enhancing the usability of quantum annealing resources for optimization purposes.
@@
The QUBO problem, illustrated by Eq. 10, can be depicted as a weighted (and/or directed) graph. In this representation, the vertices correspond to binary variables $x_{i}$, and the edges signify the quadratic terms $x_{i}x_{j}$ in the problem formulation. Linear coefficients are attributed as weights to the vertices, while quadratic coefficients are similarly assigned to the edges.
The actual D-Wave QPU topology uses a Pegasus graph of degree 15 [29]. The QUBO graph must be mapped onto that architecture, which is achieved by minor-embedding [30] i.e. mapping each logical binary variable to a group of physical qubits on the actual machine so all required connections are realized. As the QUBO function based on the generalized modularity matrix (see Eq. 10) is dense, we used complete graph embedding predefined routine (clique embedding). To reduce the time of the embedding procedure, we used the cache mechanism to compute embedding for all necessary clique sizes only once and reuse it while performing the hierarchical approach described by Algorithm 1.
@@
To compare the performance of the hierarchical process and pure quantum annealing, we evaluated results using four different approaches: the Louvain, Leiden, Bayan, and Hierarchical Gurobi algorithms, as well as a hybrid quantum-classical method based on discrete quadratic optimization models (DQM) [10, 12, 14, 25]. For the Leiden and Louvain algorithms, we computed multiple instances for each network. Specifically, each algorithm was run a certain number of times per network, and the community structure with the highest modularity was retained as the “true” structure. The DQM method demonstrated high stability, likely due to its problem partitioning strategy for fitting into the annealer [24, 25]. The Hierarchical Gurobi used the same algorithms as the Hierarchical annealing but only used the classical Gurobi solver. For the Bayan, which relies heavily on the Gurobi optimizer and its corresponding license, the Louvain, and Leiden algorithms we ran 20 instances per network. The Bayan algorithm is known to return exact solutions for networks up to a considerable size [15].
Community detection algorithms often yield varying structures due to the use of heuristics or inherent solution degeneracy [31]. We evaluated the variability between communities identified by two algorithms through 1) consensus classification matrices and 2) the Dice-Sørensen score. For each network, an $N\times N$ consensus coclassification matrix recorded how often nodes $i$ and $j$ were grouped into the same community across algorithm runs [32]. Perfect agreement between algorithms would result in matrix entries $M_{ij}$ of 0 or 1 (or 0% and 100%), indicating stable or well-defined community structures. Conversely, a range of intermediate $M_{ij}$ values suggests inconsistent community assignments. While informative, this matrix does not assess community structure optimality, necessitating the additional use of a modularity index. To compare community structures directly, we computed the Dice score between pairs of communities as:
@@
where $Z$ is in place for any of the classical alternatives described.
@@
We generated random networks using three different models. First, the Erdős-Rényi model [33] connected any pair of initially disconnected nodes with a constant probability $p$, producing random networks without clear statistical organization. Despite this, non-trivial geometrical behaviors can still emerge, making it a useful null model for study. Second, the Barabasi-Albert model [34] added new nodes dynamically, where the probability of a new node $i$ connecting to node $j$ was proportional to $j$’s degree, $p_{j}\propto k_{j}$. This preferential attachment mechanism produced networks with robust scaling properties and enhanced resilience. Finally, we used the Holme-Kim model [35], which incorporates preferential attachment but adds a triad formation step to boost clustering, defined as the presence of connected triangles in the network. This produced highly clustered, power-law networks. To test our approach in directed networks, we also used the preferential attachment mechanism applied to directed versions of the graphs [36]. This yielded power-law and scale-free networks with directed edges. The parameters of this model were left as the default values, both in Bollobás, et al., 2003 [36], and in the NetworkX Python package [37].
For all of these networks, we fitted a discrete power-law distribution to their node degrees present in a given network [38]. A network is said to have a power-law distribution or structure if the probability $P(k)$ of a given node to have degree $k$ follows an exponential distribution,